If the coefficient of real expansion is 1% more than the coefficient of apparent expansion find linear expansion?

Grade:11

1 Answers

Arun
6 years ago
The coefficient of real expansion equals the sum of coeffiient of apparent expansion and the coefficient of expansion of vessel.(all being volume expansion. 
Ar=Aa+Ag
Ag=3*Lg(coefficient of linear expansion)
Ar - Aa = Ag
1 = 3 Lg
Lg = 1/3 = 0.333...%
Regards
Arun (askIITians forum expert)